What mass of nitric 

oxide can be made from the reaction of 8.00 g NH3 with 17.0 g O2?

4NH3(g) + 5O2(g)  4NO(g) + 6H2O


1
Expert's answer
2021-10-22T03:52:01-0400

v(NH3)=8/17=0.47mol

v(O2)=17/32=0.53mol

v(NO)=4/5(O2)=0.425mol

m(NO)=0.425*30=12.75g